A stock standard solution of glucose containing 1 g of glucose/5,000mL. The working standard is prepared by diluting 5mL of the stock solution to 100mL. What is the concentration of glucose in the working standard in mg/dL?
Answer #1

A stock solution contain 1 g glucose per 5000 ml.

Hence, concentration of stock glucose solution = 1 g / 5000 ml = 2 \times 10 -04 g / ml.

Working standard is prepared by diluting 5 ml of the stock solution to 100 ml. To find out concentration of dilute solution , we can use dilution formula.

We have dilution formula, C stock \times V stock = C dilute \times V dilute

We have, C stock = 2 \times 10 -04 g / ml , V stock = 5 ml & V dilute = 100 ml.

C dilute = C stock \times V stock /  V dilute

C dilute = (2 \times 10 -04 g / ml ) \times 5 ml / 100 ml = 1 \times 10 -05 g / ml

Concentration of glucose in working standard = 1 \times 10 -05 g / ml.

We have relation, 1 g = 1000 mg & 1 dL = 100 ml. Therefore, 1 ml = 0.01 dL.

Concentration of glucose working standard = 1 \times 10 -05 g \times   ( 1000 mg / 1 g ) / ( 1 ml \times ( 1 dL / 100 ml ) )

Concentration of glucose working standard = 0.01 mg / 0.01 dL = 1 mg / dL

ANSWER : Concentration of Glucose in working standard = 1 mg / dL